# Minimalist Explanation Generation and Circuit Discovery

### v1

@@ Abstract
Machine learning models, by virtue of training, learn a large repertoire of decision rules for any given input, and any one of these may suffice to justify a prediction. However, in high-dimensional input spaces, such rules are difficult to identify and interpret. In this paper, we introduce an activation-matching–based approach to generate minimal and faithful explanations for the decisions of pre-trained image classifiers. We aim to identify minimal explanations that not only preserve the model’s decision but are also concise and human-readable. To achieve this, we train a lightweight autoencoder to produce binary masks that learns to highlight the decision-wise critical regions of an image while discarding irrelevant background. The training objective integrates activation alignment across multiple layers, consistency at the output label, priors that encourage sparsity, and compactness, along with a robustness constraint that enforces faithfulness. The minimal explanations so generated also lead us to mechanistically interpreting the model internals. In this regard we also introduce a circuit readout procedure wherein using the explanation’s forward pass and gradients, we identify active channels and construct a channel-level graph, scoring inter-layer edges by ingress weight magnitude times source activation and feature-to-class links by classifier weight magnitude times feature activation. Together, these contributions provide a practical bridge between minimal input-level explanations and a mechanistic understanding of the internal computations driving model decisions.
@@ 1 Introduction
In this work, we propose an activation-matching based approach that, given an input image and a frozen pretrained classifier, trains a lightweight autoencoder to generate a binary mask selecting the minimal set of pixels needed to preserve the model’s prediction and internal activations. We further leverage these explanations to uncover concise, channel-level views of the model’s computation, revealing sparse, data-dependent subcircuits sufficient for the decision. Beyond the forward activation pathways, we also couple this analysis with gradient information, tracing back the most prominent gradients from the output class logit toward earlier layers and the input. This dual perspective—forward activations and backward gradients—enables us to connect minimal input-level explanations with both mechanistic insight into signal flow and attribution of decision-critical pathways, yielding a more comprehensive understanding of how deep networks arrive at their predictions.
@@ 2.1 Model Explainability via Inversion
Inversion techniques aim to reconstruct input patterns that trigger specific outputs or internal activations in a neural network. Unlike explanations, which are inherently tied to a particular input and its corresponding decision, inversion seeks to synthesize representative stimuli that reveal what a model has learned in aggregate. Early work on multilayer perceptrons applied gradient-based inversion to visualize decision boundaries, though the resulting reconstructions were often noisy or adversarial-like Kindermann & Linden (1990); Jensen et al. (1999); Saad & Wunsch (2007). To address these limitations, researchers explored evolutionary search and constrained optimization Wong (2017). Subsequent advances incorporated prior-based regularization, such as smoothness constraints or pretrained generative models, which enhanced both realism and interpretability of reconstructions Mahendran & Vedaldi (2014); Yosinski et al. (2015); Mordvintsev et al. (2015); Nguyen et al. (2016; 2017). More recently, methods have emerged that stabilize inversion by learning surrogate loss landscapes Liu et al. (2022), while generative approaches conditionally reconstruct inputs likely to produce a given output Suhail & Sethi (2024). Alternative formulations as in  Suhail (2024) encode classifiers into CNF constraints, framing inversion as a deterministic satisfiability problem.
@@ 2.2 Input-Level Explainability
While inversion focuses on global characterizations of model behavior, input-level explanation methods aim to provide faithful rationales for specific predictions. Explainable AI has thus developed into a central research field Ali et al. (2023); Hsieh et al. (2024); Gilpin et al. (2018), driven by the demand for trust, transparency, and accountability in high-stakes applications. Among post-hoc attribution methods, LIME builds local surrogate models to approximate decision boundaries Hamilton et al. (2022), whereas Grad-CAM highlights salient image regions through gradient-weighted activations Selvaraju et al. (2019). More recent advances emphasize concept-based explanations that map predictions to semantically interpretable parts Lee et al. (2025). Evaluating such methods remains an open challenge: surveys have highlighted the importance of rigorous metrics combining fidelity, stability, and human-centered evaluation Zhou et al. (2021). Explanations are also being embedded into interactive systems, allowing users to guide, debug, or refine models through explanation-driven feedback Teso et al. (2022). Beyond heuristics,  Ignatiev et al. (2018) also explore abductive reasoning approaches that provide subset- or cardinality-minimal explanations with formal guarantees.
@@ 2.3 Mechanistic Interpretability of Circuits
Mechanistic interpretability investigates the circuits within a model—sparse subnetworks of neurons and connections that implement particular algorithms. Minimal explanations can reveal the smallest sufficient evidence for a prediction, offering insights into how internal components drive decisions. Early studies of circuits relied on manual inspection, limiting scalability. Recent approaches automate this process: ACDC Conmy et al. (2023) introduced a systematic framework that rediscovered known transformer circuits through activation patching. Building on this,  Rajaram et al. (2024) extended circuit discovery to vision models, extracting subnetworks responsible for concept recognition and demonstrating that targeted edits could alter predictions and enhance robustness. Further work Nainani et al. (2024) explored how circuits generalize across diverse inputs, revealing that networks often reuse a shared set of components while flexibly adapting their connectivity—a manifestation of representational superposition.
@@ 3.1 Explanation Generation
In order to generate explanations we train a lightweight autoencoder to produce a binary mask $m$, defining the explanation as $e=m\odot x,$ which suppresses irrelevant regions as proposed in Suhail et al. (2025). The autoencoder is optimized with a composite objective whose terms are weighted to balance fidelity, sparsity, smoothness, and robustness.
@@ 6 Comparisons
Figure 4: Left: Comparison with Grad-CAM. Right: Comparison with Attention Maps.
@@ 6.1 Comparisons with Grad-CAM
We begin by comparing our explanations against Grad-CAM visualizations across a variety of models and input images in the left panel of Figure 4. The first example is an image classified as house finch by EfficientNet with a confidence of 0.11. Using our method with loss weights $\lambda_{\text{act}}=1.0$, $\lambda_{\text{CE}}=4.0$, $\lambda_{\text{KL}}=0.4$, $\lambda_{\text{area}}=15.0$, $\lambda_{\text{bin}}=0.3$, $\lambda_{\text{tv}}=15.0$, and $\lambda_{\text{rob}}=6.0$, we obtain an explanation that sharply highlights the bird itself. In contrast, Grad-CAM focuses on the beak and background, producing a more diffused attribution.
@@ 6.2 Comparisons with Attention Maps
We next compare our explanations with attention maps extracted from ViT-16 shown on the right of Figure 4. The first example is an image classified as basketball with confidence 0.89. Using our loss weights $\lambda_{\text{act}}=4.5$, $\lambda_{\text{CE}}=15.0$, $\lambda_{\text{KL}}=5.0$, $\lambda_{\text{area}}=55.0$, $\lambda_{\text{bin}}=1.5$, $\lambda_{\text{tv}}=57.0$, and $\lambda_{\text{rob}}=45.0$, the explanation minimally highlights only the ring that determines the classification, yielding a confidence of 0.90. Attention maps, however, emphasize broader regions dominated by red color patches, diluting the evidence.
@@ 6.3 Comparisons with Abductive Explanations
We also compare against abductive explanations using a custom three-layer CNN trained on MNIST digits as shown in Figure  5). In this setup, the abductive explanations highlight in red and blue the minimal evidence required for the prediction, shown on the left of each example. Our method, shown on the right, instead identifies the minimal set of pixels that not only preserves the predicted class label but also maintains the classifier’s confidence relative to the original input.
@@ 7 Conclusion
We presented an activation-matching–based framework for generating minimal and faithful explanations of pre-trained image classifiers using an autoencoder that learns to produce binary masks discarding irrelevant pixels in the explanations. Beyond input-level interpretability, we further showed how these explanations can be coupled with forward activations and backward gradients to uncover sparse computational sub-circuits that realize individual decisions within deep networks. As future work, a natural direction is to explore formal guarantees of minimality in the generated explanations, strengthening the theoretical foundation of our approach while extending its applicability to broader domains and architectures.

### v2

# EXP-CAM: Explanation Generation and Circuit Discovery Using Classifier Activation Matching

@@ Abstract
Machine learning models, by virtue of training, learn a large repertoire of decision rules for any given input, and any one of these may suffice to justify a prediction. However, in high-dimensional input spaces, such rules are difficult to identify and interpret. In this paper, we introduce EXP-CAM: an explanation generation and circuit discovery approach using Classifier Activation Matching. EXP-CAM can generate minimal and faithful explanations for the decisions of pre-trained image classifiers that not only preserve the model’s decision but are also concise and human-readable. We aim to identify minimal explanations that not only preserve the model’s decision but are also concise and human-readable. To achieve this, we train a lightweight autoencoder to produce binary masks that learns to highlight the decision-wise critical regions of an image while discarding irrelevant background. The training objective integrates activation alignment across multiple layers, consistency at the output label, priors that encourage sparsity, and compactness, along with a robustness constraint that enforces faithfulness. The minimal explanations so generated also lead us to mechanistically interpreting the model internals. In this regard we also introduce a circuit readout procedure wherein using the explanation’s forward pass and gradients, we identify active channels and construct a channel-level graph, scoring inter-layer edges by ingress weight magnitude times source activation and feature-to-class links by classifier weight magnitude times feature activation. Together, these contributions provide a practical bridge between minimal input-level explanations and a mechanistic understanding of the internal computations driving model decisions.
@@ 1 Introduction
In this work, we propose EXP-CAM as an Activation-Matching based approach that, given an input image and a frozen pretrained classifier, trains a lightweight autoencoder to generate a binary mask selecting the minimal set of pixels needed to preserve the model’s prediction and internal activations. We further leverage these explanations to uncover concise, channel-level views of the model’s computation, revealing sparse, data-dependent subcircuits sufficient for the decision. Beyond the forward activation pathways, we also couple this analysis with gradient information, tracing back the most prominent gradients from the output class logit toward earlier layers and the input. This dual perspective—forward activations and backward gradients—enables us to connect minimal input-level explanations with both mechanistic insight into signal flow and attribution of decision-critical pathways, yielding a more comprehensive understanding of how deep networks arrive at their predictions.
Our main contributions are as follows:
  1. 1.
    We present an explanation generation framework that produces minimal explanations guaranteed to replicate the model’s behavior by aligning internal activations and output fidelity.
  2. 2.
    We introduce a sparse causal circuit discovery process corresponding to these minimal explanations, identifying the dominant activation pathways within the network.
  3. 3.
    EXP-CAM only requires per image training of the mask generator from scratch—for both seen and unseen samples—within seconds.
  4. 4.
    While methods like Grad-CAM yield a single fixed explanation, our EXP-CAM framework supports generating multiple explanations with controllable levels of minimality and compactness through adjustable mask priors.
@@ 2.1 Model Explainability via Inversion
Inversion techniques aim to reconstruct input patterns that trigger specific outputs or internal activations in a neural network. Unlike explanations, which are inherently tied to a particular input and its corresponding decision, inversion seeks to synthesize representative stimuli that reveal what a model has learned in aggregate. Early work on multilayer perceptrons applied gradient-based inversion to visualize decision boundaries, though the resulting reconstructions were often noisy or adversarial-like KINDERMANN1990277; 784232; SAAD200778. To address these limitations, researchers explored evolutionary search and constrained optimization Wong2017NeuralNI. Subsequent advances incorporated prior-based regularization, such as smoothness constraints or pretrained generative models, which enhanced both realism and interpretability of reconstructions mahendran2015understanding; yosinski2015understanding; mordvintsev2015inceptionism; nguyen2016synthesizing; nguyen2017plug. More recently, methods have emerged that stabilize inversion by learning surrogate loss landscapes liu2022landscapelearningneuralnetwork, while generative approaches conditionally reconstruct inputs likely to produce a given output suhail2024networkcnn. Alternative formulations as in  suhail2024network encode classifiers into CNF constraints, framing inversion as a deterministic satisfiability problem.
@@ 2.2 Input-Level Explainability
While inversion focuses on global characterizations of model behavior, input-level explanation methods aim to provide faithful rationales for specific predictions. Explainable AI has thus developed into a central research field ALI2023101805; hsieh2024comprehensiveguideexplainableai; Gilpin2018ExplainingEA, driven by the demand for trust, transparency, and accountability in high-stakes applications. Among post-hoc attribution methods, LIME builds local surrogate models to approximate decision boundaries 10022096, whereas Grad-CAM highlights salient image regions through gradient-weighted activations Selvaraju_2019. More recent advances emphasize concept-based explanations that map predictions to semantically interpretable parts 10.1007/978-3-031-92648-8_17. Evaluating such methods remains an open challenge: surveys have highlighted the importance of rigorous metrics combining fidelity, stability, and human-centered evaluation electronics10050593.
Explanations are also being embedded into interactive systems, allowing users to guide, debug, or refine models through explanation-driven feedback teso2022leveragingexplanationsinteractivemachine. Beyond heuristics,  ignatiev2018abductionbasedexplanationsmachinelearning also explore abductive reasoning approaches that provide subset- or cardinality-minimal explanations with formal guarantees.
@@ 2.3 Mechanistic Interpretability of Circuits
Mechanistic interpretability investigates the circuits within a model—sparse subnetworks of neurons and connections that implement particular algorithms. Minimal explanations can reveal the smallest sufficient evidence for a prediction, offering insights into how internal components drive decisions. Early studies of circuits relied on manual inspection, limiting scalability. Recent approaches automate this process: ACDC conmy2023automatedcircuitdiscoverymechanistic introduced a systematic framework that rediscovered known transformer circuits through activation patching. Building on this,  rajaram2024automaticdiscoveryvisualcircuits extended circuit discovery to vision models, extracting subnetworks responsible for concept recognition and demonstrating that targeted edits could alter predictions and enhance robustness. Further work nainani2024adaptivecircuitbehaviorgeneralization explored how circuits generalize across diverse inputs, revealing that networks often reuse a shared set of components while flexibly adapting their connectivity—a manifestation of representational superposition.
@@ 3.1 Explanation Generation
In order to generate explanations we train a lightweight autoencoder to produce a binary mask $m$, defining the explanation as $e=m\odot x,$ which suppresses irrelevant regions as proposed in suhail2025activationmatchingexplanationgeneration. The autoencoder is optimized with a composite objective whose terms are weighted to balance fidelity, sparsity, smoothness, and robustness.
@@ 6 Comparisons
Figure 4: Comparison with Grad-CAM.
@@ 6.1 Comparisons with Grad-CAM
We begin by comparing our explanations against Grad-CAM visualizations across a variety of models and input images in the Figure 4. The first example is an image classified as house finch by EfficientNet with a confidence of 0.11. Using our method with loss weights $\lambda_{\text{act}}=1.0$, $\lambda_{\text{CE}}=4.0$, $\lambda_{\text{KL}}=0.4$, $\lambda_{\text{area}}=15.0$, $\lambda_{\text{bin}}=0.3$, $\lambda_{\text{tv}}=15.0$, and $\lambda_{\text{rob}}=6.0$, we obtain an explanation that sharply highlights the bird itself. In contrast, Grad-CAM focuses on the beak and background, producing a more diffused attribution.
@@ 6.2 Comparisons with Attention Maps
We next compare our explanations with attention maps extracted from ViT-16 shown in the Figure 5. The first example is an image classified as basketball with confidence 0.89. Using our loss weights $\lambda_{\text{act}}=4.5$, $\lambda_{\text{CE}}=15.0$, $\lambda_{\text{KL}}=5.0$, $\lambda_{\text{area}}=55.0$, $\lambda_{\text{bin}}=1.5$, $\lambda_{\text{tv}}=57.0$, and $\lambda_{\text{rob}}=45.0$, the explanation minimally highlights only the ring that determines the classification, yielding a confidence of 0.90. Attention maps, however, emphasize broader regions dominated by red color patches, diluting the evidence.
@@ 6.3 Comparisons with Abductive Explanations
We also compare against abductive explanations using a custom three-layer CNN trained on MNIST digits as shown in Figure  6). In this setup, the abductive explanations highlight in red and blue the minimal evidence required for the prediction, shown on the left of each example. Our method, shown on the right, instead identifies the minimal set of pixels that not only preserves the predicted class label but also maintains the classifier’s confidence relative to the original input.
@@ 7 Conclusion
We presented EXP-CAM as an activation-matching–based framework for generating minimal and faithful explanations of pre-trained image classifiers using an autoencoder that learns to produce binary masks discarding irrelevant pixels in the explanations. Beyond input-level interpretability, we further showed how these explanations can be coupled with forward activations and backward gradients to uncover sparse computational sub-circuits that realize individual decisions within deep networks. As future work, a natural direction is to explore formal guarantees of minimality in the generated explanations, strengthening the theoretical foundation of our approach while extending its applicability to broader domains and architectures.